# A GPU-accelerated Large-scale Simulator for Transportation System Optimization Benchmarking

### v1

@@ 3.2 Controllable Objects
Traffic Signal. The simulator allows the user to set the traffic signal control policies for given junctions via engine.set_tl_policy(id, policy). The policy enumeration includes MANUAL, FIXED_TIME, MAX_PRESSURE and NONE. Under the MANUAL policy, the user can change the current phase and duration of the signal via engine.set_tl_phase(id, phase_index) and engine.set_tl_duration(id, duration). The FIXED_TIME policy indicates that the fixed phase procedure built into the map data is used. The MAX_PRESSURE policy indicates that the adaptive maximum pressure algorithm [31] is used. The NONE policy indicates that there is no signaling.
Lane. Lanes in the simulator include both clearly marked lanes on the roadway and "virtual" lanes within junctions that connect the two roadways. For lanes, the user can first set their maximum speed via engine.set_lane_max_speed(id, max_speed). Secondly, the user can set whether the lane is restricted from passing via engine.set_lane_restriction(id, flag).
@@
Vehicle. The user can change the route of the vehicle via engine.set_vehicle_route (vehicle_id, route, end_lane, end_s) to modify its route and destination lane position.
@@ 5 Experiments
Simulator Performance. To illustrate the computational performance of our proposed simulators, we compared the computational efficiency of simulators including SUMO, CityFlow, CBLab, and our proposed simulator for different road network scales and vehicle sizes.
@@
Figure 3: The performance comparison with the number of vehicles. (best view in color)
@@
We adopt the datasets from CBLab [19] with the MIT License, which includes 6 real-world city datasets and 9 synthesized datasets. We simulated 3600 steps for all the datasets and record the total running times as the performance of each simulator. All simulations are conducted in the same hardware environment with an Intel(R) Xeon(R) Platinum 8462Y CPU (64 threads) and an NVIDIA GeForce RTX 4090 GPU. As shown in Figure 3, the result indicates that our proposed simulator has a huge performance improvement over existing simulators. On the largest dataset, the running time of ours is 42.81s and that of the best baseline (CityFlow) is 3806.7s, a relative performance improvement of 88.09 times.
@@
To benchmark the optimization algorithms for the five transportation system optimization scenarios described above, we chose Beijing, Shanghai, Paris and New York as test cities. The road networks of these cities are built using our mosstool toolchain. The real origin-destination (OD) matrices of these cities are also generated by mosstool using generative AI methods. As synthetic datasets, in terms of vehicle departure times, we kept only the morning and evening peaks to challenge the optimization algorithms for each scenario. The total number of vehicles was scaled based on the generated real OD matrix to construct travel demand data for three different congestion levels including smooth (marked as City-S), normal (marked as City-N), congested (marked as City-C). More information on the synthetic datasets will be provided in Appendix A. We evaluate the optimization effectiveness of different algorithms under the above cities and congestion levels, using ATT and TP as global metrics for comparison. In the following text, the comparisons of the various optimization algorithms used in all the five scenarios and their performance under normal congestion will be reported. The detailed experimental settings and the complete results are presented in Appendix B and Appendix C respectively due to the page limit.
@@
Traffic Signal Control Benchmark. In this scenario, the task is to choose the best traffic signal signal phase from the list of available phases for each junction. We compared the rule-based algorithms, including fixed-time algorithm [15] and maximum pressure algorithm [31], and the reinforcement learning-based algorithms, including FRAP [42], MPLight [3], CoLight [35], Efficient-MPLight [38], Advanced-MPLight and Advanced-Colight [40], as well as a pressure-based model trained with PPO [27]. The related algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 2.
@@
Table 2: The benchmark results for the traffic signal control scenario with normal traffic conditions.
@@
FixedTime 4843.05 120049 4324.07 131020 4245.86 60020 4682.07 72725
MaxPressure 4580.41 132055 4045.75 144640 3984.20 64405 4309.91 83927
FRAP 5105.22 112321 4671.52 121896 4404.23 58481 5002.48 68196
MPLight 4790.12 124991 4674.93 122198 3980.07 64921 4196.37 85331
CoLight 5108.88 112672 4640.50 124565 4413.91 58513 4989.17 68184
Efficient-MPLight 5101.14 113272 4480.28 132995 4364.89 60428 5019.92 67303
Advanced-MPLight 5049.47 117568 4603.57 127911 4281.72 61470 5031.15 67322
Advanced-CoLight 5107.23 112778 4661.07 123336 4408.14 58929 5014.66 68292
PPO 4452.05 136630 4143.19 141768 4017.61 64277 4254.92 84411
Dynamic Lane Assignment within Junctions Benchmark. In this scenario, the task is to assign the direction, e.g. left or straight, for the in-going lanes of each junction. We compared the following methods: 1) NoChange, where we do not change the direction of the lane and leave it as it is, 2) Random, where we randomly change the direction in every period, 3) Rule, where we estimate the number of vehicles going for each direction and choosing the direction with the maximum number of vehicles, 4) PPO, where we use a PPO-trained model to estimate the number of vehicles. The above algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 3.
Table 3: The benchmark results for the dynamic lane assignment scenario with normal traffic conditions.
@@
NoChange 4846.70 119890 4322.21 131145 4245.84 60020 4674.08 72870
Random 4839.71 120338 4324.96 131216 4176.70 61810 4636.11 74055
Rule 4761.33 123673 4258.22 133346 4155.11 62366 4615.01 74254
PPO 4769.98 122929 4256.51 133379 4160.89 61792 4614.52 73907
@@
Table 4: The benchmark results for the tidal lane control scenario with normal traffic conditions.
@@
NoChange 4844.57 120105 4334.80 130604 4224.94 60794 4675.10 72834
Random 4827.75 120778 4338.40 130283 4216.17 60946 4665.83 73335
Rule 4823.40 120901 4313.48 131315 4192.85 61636 4638.03 74284
PPO 4820.48 120936 4304.29 132167 4187.37 61738 4628.47 74756
Congestion Pricing Benchmark. In this scenario, each driver has three candidate routes and the task is to set the price of each road to motivate drivers to choose the route that avoids congested areas. We compared $\Delta$-toll [28] and EBGtoll [26] with two baselines: 1) NoChange, where we do not set the prices, 2) Random, where the drivers randomly choose a route. The above algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 5.
Table 5: The benchmark results for the congestion pricing scenario with normal traffic conditions.
@@
NoChange 4840.18 120207 4328.34 130621 4239.23 60100 4681.75 72765
Random 5190.37 105640 4422.00 131346 4144.47 62284 4830.86 68976
$\Delta$-toll 4667.77 131747 4096.62 147533 4040.68 65024 4549.26 78182
EBGtoll 5637.30 80476 4637.64 116624 4240.82 60362 5096.60 59154
Road Planning Benchmark. In this scenario, the algorithms are asked to select at most 30 roads from 50 candidates for construction to minimize post-construction ATT. We compared 5 methods: 1) NoChange, no of these 50 roads are built, 2) Random, where we select random roads to build, 3) Rule-based, where we select the top-30 vehicle count roads to build, 4) simulated annealing [14], 5) bayesian optimization [6]. The above algorithms are tested both on morning peak from 6:00 to 12:00 and evening peak from 17:00 to 23:00 and computes the mean of metrics. The result are presented in Table 6.
Table 6: The benchmark results for the road planning scenario with normal traffic conditions.
@@
No-Change 8439.44 161722 6699.66 161722 7193.90 76327 7892.50 105533
Random 8304.71 163148 6567.12 181063 7106.19 75839 7967.24 102996
Rule 8235.49 164247 6570.72 181504 7177.88 76176 7956.08 102951
SA 8332.44 163660 6590.66 180954 7154.16 76164 7871.23 105507
GeneralBO 8242.19 164182 6721.78 178790 7161.74 75979 7759.88 106883
@@ 6 Conclusion
In this paper, we propose a high-performance large-scale microscopic simulator powered by GPU for transportation system simulation and optimization. We also benchmarked the effect of different optimization algorithms on five transportation system optimization scenarios with different traffic flows in four cities. Interested researchers can use the same pipeline to benchmark most cities around the world with our open source simulator and toolchain. We believe that the proposed simulator will contribute to more researchers joining the research work on urban transportation system optimization. We hope that this will not only support more research work on transportation system optimization scenarios, but also promote the development of urban transportation systems towards AI-driven intelligent transportation systems.
Checklist
@@
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [No]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] We do not think that the traffic simulator will have any negative societal impacts.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
@@
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A] There is no theoretical result in the paper.
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A] There is no theoretical result in the paper.
@@
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Provided in our codes URL.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Provided in Appendix A and Appendix B.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [No]
@@
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We use CBLab datasets for performance evaluation and already cite it in the paper.
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] We provide all the scenario codes and our synthetic datasets in the Github URL.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] The data are generated from public data sources.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] The data are generated without any personally identifiable information or offensive content.
@@
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A] No crowdsourcing was used and no human subjects were included.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A] No crowdsourcing was used and no human subjects were included.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A] No crowdsourcing was used and no human subjects were included.
Supplementary material
We present the following items in the supplementary material section:
  1. 1.
    Datasets for transportation system optimization benchmarking. (Section A)
  2. 2.
    The settings of transportation system optimization scenarios. (Section B)
  3. 3.
    Complete benchmark results. (Section C)
Appendix A Datasets for Transportation System Optimization Benchmarking

### v2

@@ 3.2 Controllable Objects
Traffic Signal. The simulator allows the user to set the traffic signal control policies for given junctions via engine.set_tl_policy(id, policy). The policy enumeration includes MANUAL, FIXED_TIME, MAX_PRESSURE and NONE. Under the MANUAL policy, the user can change the current phase and duration of the signal via engine.set_tl_phase(id, phase_index) and engine.set_tl_duration(id, duration). The FIXED_TIME policy indicates that the fixed phase procedure built into the map data is used. The MAX_PRESSURE policy indicates that the adaptive maximum pressure algorithm (Varaiya, 2013) is used. The NONE policy indicates that there is no signaling.
Lane. Lanes in the simulator include both clearly marked lanes on the roadway and ”virtual” lanes within junctions that connect the two roadways. For lanes, the user can first set their maximum speed via engine.set_lane_max_speed(id, max_speed). Secondly, the user can set whether the lane is restricted from passing via engine.set_lane_restriction(id, flag).
@@
Vehicle. The user can change the route of the vehicle via engine.set_vehicle_route (vehicle_id, route, end_lane, end_s) to modify its route and destination.
@@ 5 Experiments
((a))
((b))
@@
Figure 3: (a) The performance comparison with different sizes of road networks and number of vehicles. (b) Comparison of real-world and simulated average vehicle speeds. (best view in color)
@@
Simulator Performance. To illustrate the computational performance of our proposed simulators, we compared the computational efficiency of simulators including SUMO, CityFlow, CBLab, and our proposed simulator for different road network scales and vehicle sizes. We adopt the datasets from CBLab (Liang et al., 2023) and choose three road networks of different sizes. For each road networks, we choose three travel demands of different sizes. We simulated 3600 steps for all the datasets and record the total running times as the performance of each simulator. As shown in Figure 3(a), the result indicates that our proposed simulator has a huge performance improvement over existing simulators under all conditions. On the largest dataset, the running time of ours is 42.81s and that of the best baseline (CityFlow) is 3806.7s, a relative performance improvement of 88.09 times. The statistics of the datasets and hardware used are left in Appendix C.
Simulator Realism. To indicate its realism, we compared the similarity between our and CityFlow’s (the best baseline in performance comparison) simulation results and the real traffic situation dataset from Shenzhen, China (Yu et al., 2023). The dataset contains 4.22 million vehicle GPS trajectories as the ground truth of average vehicle speeds and 156,856 vehicle trips captured by traffic cameras as the input. The experiment compares the simulated average vehicle speeds of 1,341 roads from 8 to 9 AM and the results are shown in Figure 3(b). The average road speeds obtained from the our simulator (RMSE=8.5km/h, correlation coefficient=0.7691) are closer to the real-world data compared to CityFlow (RMSE=16km/h, correlation coefficient=0.5286), which shows that our simulator achieves a more realistic simulation by adopting the newer car-following model (IDM) and lane-change model (randomized MOBIL). The visualization results of road traffic status are shown in Figure A3.
@@
To benchmark the optimization algorithms for the five scenarios described above, we chose Beijing, Shanghai, Paris and New York as test cities. The road networks of these cities are built using the toolchain. The real origin-destination (OD) matrices of these cities are also generated by the toolchain using generative AI methods. As synthetic datasets, in terms of vehicle departure times, we kept only the morning and evening peaks to challenge the optimization algorithms for each scenario. The total number of vehicles was scaled based on the generated real OD matrix to construct travel demand data for three different congestion levels including smooth (marked as City-S), normal (marked as City-N), congested (marked as City-C). More information on the synthetic datasets will be provided in Appendix E. We evaluate the optimization effectiveness of different algorithms under the above cities and congestion levels, using ATT and TP as global metrics for comparison. In the following text, the comparisons of the various optimization algorithms used in all the five scenarios and their performance under normal congestion will be reported. The detailed experimental settings and the complete results are presented in Appendix F and Appendix G, respectively.
@@
Traffic Signal Control Benchmark. In this scenario, the task is to choose the best traffic signal signal phase from the list of available phases for each junction. We compared the rule-based algorithms, including fixed-time algorithm (Koonce & Rodegerdts, 2008) and maximum pressure algorithm (Varaiya, 2013), and the reinforcement learning-based algorithms, including FRAP (Zheng et al., 2019), MPLight (Chen et al., 2020), CoLight (Wei et al., 2019), Efficient-MPLight (Wu et al., 2021), Advanced-MPLight and Advanced-Colight (Zhang et al., 2022), as well as a pressure-based model trained with PPO (Schulman et al., 2017). The related algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 2.
@@
Table 2: The results for the traffic signal control scenario with normal traffic conditions.
@@
FixedTime 4,843.1 120,049 4,324.1 131,020 4,245.9 60,020 4,682.1 72,725
MaxPressure 4,580.4 132,055 4,045.8 144,640 3,984.2 64,405 4,309.9 83,927
FRAP 5,105.2 112,321 4,671.5 121,896 4,404.2 58,481 5,002.5 68,196
MPLight 4,790.1 124,991 4,674.9 122,198 3,980.1 64,921 4,196.4 85,331
CoLight 5,108.9 112,672 4,640.5 124,565 4,413.9 58,513 4,989.2 68,184
Efficient-MPLight 5,101.1 113,272 4,480.3 132,995 4,364.9 60,428 5,019.9 67,303
Advanced-MPLight 5,049.5 117,568 4,603.6 127,911 4,281.7 61,470 5,031.1 67,322
Advanced-CoLight 5,107.2 112,778 4,661.1 123,336 4,408.1 58,929 5,014.7 68,292
PPO 4,452.1 136,630 4,143.2 141,768 4,017.6 64,277 4,254.9 84,411
Dynamic Lane Assignment within Junctions Benchmark. In this scenario, the task is to assign the direction, e.g. left or straight, for the in-going lanes of each junction. We compared the following methods: 1) NoChange, where we keep lane direction unchanged, 2) Random, where we randomly change the direction in every period, 3) Rule, where we estimate the number of vehicles going for each direction and choosing the direction with the maximum number of vehicles, 4) PPO, where we use a PPO-trained model to estimate the number of vehicles. The above algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 3.
Table 3: The results for the dynamic lane assignment scenario with normal traffic conditions.
@@
NoChange 4,846.7 119,890 4,322.2 131,145 4,245.8 60,020 4,674.1 72,870
Random 4,839.7 120,338 4,325.0 131,216 4,176.7 61,810 4,636.1 74,055
Rule 4,761.3 123,673 4,258.2 133,346 4,155.1 62,366 4,615.0 74,254
PPO 4,770.0 122,929 4,256.5 133,379 4,160.9 61,792 4,614.5 73,907
@@
Table 4: The results for the tidal lane control scenario with normal traffic conditions.
@@
NoChange 4,844.6 120,105 4,334.8 130,604 4,224.9 60,794 4,675.1 72,834
Random 4,827.8 120,778 4,338.4 130,283 4,216.2 60,946 4,665.8 73,335
Rule 4,823.4 120,901 4,313.5 131,315 4,192.9 61,636 4,638.0 74,284
PPO 4,820.5 120,936 4,304.3 132,167 4,187.4 61,738 4,628.5 74,756
Congestion Pricing Benchmark. In this scenario, each driver has three candidate routes and the task is to set the price of each road to motivate drivers to choose the route that avoids congested areas. We compared $\Delta$-toll (Sharon et al., 2017) and EBGtoll (Qiu et al., 2019) with two baselines: 1) NoChange, where we do not set the prices, 2) Random, where the drivers randomly choose a route. The above algorithms are trained and tested in the morning rush hour scenario, from 7:00 to 10:00. The results are presented in Table 5.
Table 5: The results for the congestion pricing scenario with normal traffic conditions.
@@
NoChange 4,840.2 120,207 4,328.3 130,621 4,239.2 60,100 4,681.8 72,765
Random 5,190.4 105,640 4,422.0 131,346 4,144.5 62,284 4,830.9 68,976
$\Delta$-toll 4,667.8 131,747 4,096.6 147,533 4,040.7 65,024 4,549.3 78,182
EBGtoll 5,637.3 80,476 4,637.6 116,624 4,240.8 60,362 5,096.6 59,154
Road Planning Benchmark. In this scenario, the algorithms are asked to select at most 30 roads from 50 candidates for construction to minimize post-construction ATT. We compared 5 methods: 1) NoChange, no of these 50 roads are built, 2) Random, where we select random roads to build, 3) Rule-based, where we select the top-30 vehicle count roads to build, 4) simulated annealing (SA) (Kirkpatrick et al., 1983), 5) bayesian optimization named GeneralBO (Cowen-Rivers et al., 2020). The above algorithms are tested both on morning peak from 6:00 to 12:00 and evening peak from 17:00 to 23:00 and computes the mean of metrics. The result are presented in Table 6.
Table 6: The results for the road planning scenario with normal traffic conditions.
@@
No-Change 8,439.4 161,722 6,699.7 161,722 7,193.9 7,6327 7,892.5 105,533
Random 8,304.7 163,148 6,567.1 181,063 7,106.2 75,839 7,967.2 102,996
Rule 8,235.5 164,247 6,570.7 181,504 7,177.9 76,176 7,956.1 102,951
SA 8,332.4 163,660 6,590.7 180,954 7,154.2 76,164 7,871.2 105,507
GeneralBO 8,242.2 164,182 6,721.8 178,790 7,161.7 75,979 7,759.9 106,883
@@ 6 Conclusion
In this paper, we propose a high-performance large-scale microscopic simulator powered by GPU for transportation system simulation and optimization. We also benchmarked the effect of different optimization algorithms on five transportation system optimization scenarios with different traffic flows in four cities. Interested researchers can use the same pipeline to benchmark most cities around the world with our open source simulator and toolchain. To help interested researchers quickly try out the simulator, we also build a web platform introduced in Appendix H to support no-code trials. We believe that the proposed simulator and platform will contribute to more researchers joining the research work on urban transportation system optimization. We hope that this will not only support more research work on transportation system optimization scenarios, but also promote the development of urban transportation systems towards AI-driven intelligent transportation systems.
Appendice
We present the following items in the appendix section:
@@
    Our statement on the open-source codes mentioned in the paper. (Section A)
@@
    System design, implementation details, and comparison with the baseline systems. (Section B)
@@
    Datasets and settings for simulator performance comparison. (Section C)
@@
    Visualization of simulator realism comparison. (Section D)
@@
    Datasets for transportation system optimization benchmarking. (Section E)
  6. 6.
    The settings of transportation system optimization scenarios. (Section F)
  7. 7.
    Complete benchmark results. (Section G)
  8. 8.
    The guidance of our web platform. (Section H)
Appendix A Statement of Open-source Codes
All the codes with documentation including the simulator, the toolchain, the benchmark codes are open-source and available at Github.
We are constantly developing the simulator and toolchain, so implementation details such as interface names may change with version updates. Replication of benchmark experiments should always be based on the specified version.
Appendix B System Design, Implementation, and Comparison
In the section, we present our system execution process to help the reader understand how GPU acceleration is implemented, and detail the vehicle model implementation and related considerations about regarding balancing computational performance with realism.
B.1 Preliminary
Figure A1: An road network example used for simulation. (best view in color)
To help better understand the system implementation presented in the section, we give the simulation object categories and their definitions as follows.
Lane. A lane is the polyline geospatial space in which a vehicle moves and contains center line coordinates. Vehicles always travel along the center line of the lane. There are topological connections between lanes to help the vehicle identify the next lane that can be reached at the end of the current lane travel. Lanes are divided into those belonging to roads and those belonging to junctions as shown in Figure A1. Lanes belonging to a road are parallel to each other without crossing, do not require traffic signals, and allow vehicles to change lanes. In contrast, lanes belonging to a junction may be crossed, requiring a signal and not allowing lane changes.
Road. A road is a collection of lanes that are parallel to each other without crossing.
Junction. A junction is a collection of lanes that may cross each other and contain a set of traffic signals.
AOI. As shown in Figure A1, an area of interest (AOI) is a polygonal geospatial space that has multiple connections to lanes for vehicle entering and exiting. The AOI can be used as a starting and ending location for the vehicle. Alternatively, the start and end locations of the vehicle can be also locations in the lane.
Person. Person is a generic term for vehicles, pedestrians, etc. Currently, only vehicles are supported for simulation.
B.2 Execution Process
Figure A2: The Step() execution process of the proposed simulator. (best view in color)
Same as the baselines, our proposed simulator uses the paradigm of discrete-time simulation for microscopic traffic simulation. The Step() function for executing each iteration is displayed in Figure A2. As shown in the figure, all simulation processes are performed on the GPU to avoid the PCI bus data transfer between the CPU and the GPU from becoming a system bottleneck. Based on the two-phase parallel process for read/write separation system design discussed in Section 3.1, the entire iterative process is divided into the prepare phase and the update phase for sequential execution. The CUDA global synchronization function cudaDeviceSynchronize() is executed after each phase to ensure that all tasks launched in that phase are finished. In each stage, in addition to using the CUDA kernel function (marked as foo$<$...$>$ in the figure) programming approach directly to perform SIMD calculations, we also use multiple CUDA asynchronous streams to simultaneously execute multiple sets of kernel functions. The above combination maximizes the utilization of GPU hardware power for computational acceleration.
In the prepare phase, there are three CUDA asynchronous streams. In Stream 1, we first perform the statistics for vehicle related data through atomic operations and then perform data replication from runtime to snapshot, which is the key task of the read/write separation strategy in the prepare phase. Subsequently, the linked-list based vehicle sensing indexes are updated including clearing side pointers for all nodes, removing nodes that left, reordering, and processing inserted nodes, and finally the side pointers will be rebuilt. To support the other features of the simulation, lane observations are updated so that the AOI knows whether its connection to the lane can put in a vehicle or not, and lane pressures are computed for adaptive traffic signals based on the maximum pressure method. In Stream 2, traffic signal phases at junctions are replicated to the corresponding lanes. In Stream 3, AOIs process people who completed their trip and entered in the previous step.
In the update phase, there are four CUDA asynchronous streams. In Stream 1, we perform simulation computations for all people in parallel, including executing vehicle simulation models will be presented in Section B.3, and updating vehicle motion states such as position and velocity based on the actions output from the models. Subsequently, if the lane the vehicle is in changes, the deletion and addition are stored in the buffer of the old lane and the new lane, respectively. If the vehicle reaches the end, it is processed accordingly, e.g., it is added to the AOI addition buffer. All vehicle’s motion statuses can be saved for file output if needed. In Stream 2, the junctions collect the pressures of entering and exiting lanes and updates the signal phases using the maximum pressure algorithm. Additionally, signal phase updates can also be made using fixed phases or specified via the Python API. In Stream 3, AOIs put vehicles reaching their departure times into the corresponding lanes if the connections are safe for collision by checking the lane observations. In Stream 4, we perform lane related statistics like average speed and output lane status including traffic signal status.
Besides, we implement the process of writing simulation output into the files on CPU, which becomes a system bottleneck due to both PCI bus bandwidth and disk write speed. However, in reinforcement learning applications, we usually do not need detailed information such as the position and speed of all vehicles at each moment, but rather statistics such as the length of queues at intersections. Also, the data is stored in memory instead of files. Therefore the performance experiments for our proposed simulator and baselines do not include writing simulation results to a file.
B.3 Vehicle Model Implementation Compared to the Baseline Systems
The implementation of the execution process described above focuses on illustrating how to simulate efficiently, and this subsection focuses on the vehicle simulation model used in our proposed simulator and the performance versus realism trade-offs considered. We will also discuss comparisons and differences among our vehicle simulation models and baseline systems.
In our simulator, the vehicle behavior is controlled by the car-following model, the lane-changing model, and the traffic signal strategy.
For the car-following model, we use the IDM model (Treiber et al., 2000). The formula is as follows:
$a(t)$ $=a_{max}\left[1-\left(\frac{v(t)}{v_{0}}\right)^{\delta}-\left(\frac{s^{*}(t)}{s(t)}\right)^{2}\right]$ (1)
$s^{*}(t)$ $=s_{0}+max\left(0,v(t)\cdot T+\frac{v(t)\cdot\Delta v(t)}{2\sqrt{a_{max}\cdot a_{comf}}}\right),$ (2)
where $a_{max}$ is the maximum acceleration of the vehicle, $a_{comf}$ is the comfortable deceleration, $T$ is the headway, $s_{0}$ is the minimum distance from the front vehicle. The four parameters are set in the vehicle’s profile. $v_{0}$ is the expected velocity of the vehicle, that is, the minimum value of road velocity limit and maximum vehicle velocity. $v(t)$, $s(t)$, and $a(t)$ are the velocity, the distance from the front vehicle, the acceleration at the moment $t$, respectively. $\Delta v(t)$ is the difference in velocity between the vehicle and the front one. $\delta$ is a hyper-parameter of the model, which is set $4.0$. In the implementation, the velocity and distance of the front vehicle is obtained by the vehicle sensing indexes with only one pointer operation. If there is no front vehicle in the lane, the first vehicle in the next lane will be used as a substitution.
For the lane-changing model, we use a randomized MOBIL model. The MOBIL model (Kesting et al., 2007) first calculate a utility by the following formula:
$u=(\tilde{a}_{ego}-a_{ego})+p\cdot\left((\tilde{a}_{new}-a_{new})+(\tilde{a}_{old}-a_{old})\right),$
where $p$ is the politeness factor which is set $0.1$. $\tilde{a}_{ego}$ and $a_{ego}$ denote the original acceleration of the vehicle and its new acceleration if the vehicle changes its lane, respectively. $\tilde{a}_{new}$ and $a_{new}$ denote the original acceleration of the new follower and its new acceleration if the ego vehicle changes its lane, respectively. $\tilde{a}_{old}$ and $a_{old}$ denote the original acceleration of the old follower and its new acceleration if the ego vehicle changes its lane, respectively. Following Feng et al. (2021), we introduce randomization to avoid identical lane-changing behavior of all vehicles and to break perfect symmetry. The total utility is calculated as follows:
$u_{T}=max(0,u_{L})+max(0,u_{R}),$
where $u_{L}$ and $u_{R}$ denote the utilities of changing lanes to the left and right lanes, respectively. Then a total lane change probability is computed as follows:
$p_{LC}=\left\{\begin{aligned} 0.9,&\quad u_{T}\geq 1\\ (0.9-2\times 10^{-8})u_{T},&\quad 0<u_{T}<1\\ 2\times 10^{-8},&\quad u_{T}=0\end{aligned}\right.$
The vehicle firstly determines whether to change lanes based on the probability $p_{LC}$, and secondly determines the direction of lane change based on $u_{L}$ and $u_{R}$. The all acceleration computation use the car-following model and the inputs are also obtained by the side pointer mechanism in the vehicle sensing indexes. Vehicles are prohibited from changing lanes away from the group of lanes that allow access to the next road.
For traffic signals, the vehicle always checks the status of the signal for the next lane and decelerates to a stop based on the IDM model (assuming a stationary vehicle at the end of the current lane) in a red or yellow light situation.
Table A1: Comparison of key components in vehicle simulation modeling.
Vehicle Behavior SUMO CityFlow CBLab Ours
Car-following Selectable Krauß Krauß IDM
Lane-changing Earliest Private Randomized MOBIL
Traffic Signal
Overlap and priority in junction - -
We compare the key components in vehicle simulation modeling among the baseline systems and our proposed simulator in the above Table A1. Car-following and Lane-changing are discussed at Section 2.1. Traffic signal indicates whether the vehicle is capable of adjusting its speed to the signal status before the junction, which is plain. Overlap and priority in junction means whether vehicles within the junction are considered for access prioritization to avoid potential intersecting vehicles. Both SUMO and CityFlow model overlap and priority in junction. However, based on the findings reported in the CBLab article (in Section 2.5.1) with our test results, modeling of overlap and priority in junction can easily lead to vehicle deadlocks (i.e., 3 or more vehicles waiting for each other) under large-scale simulations. Therefore, we adopt the same treatment as CBLab, dropping the modeling of overlap and priority in junction to ensure the practicability of the simulation in large-scale scenarios. In this setup, the realism of the junction simulation relies on conflict-free signal phase settings. This is a limitation of our proposed simulator.
Appendix C Datasets and Settings for Simulator Performance Comparison
Table A2: The statistics of the datasets used for simulator performance comparison.
Roadnet #Road #Junction #Vehicle
Roadnet-S 214 67 {19; 194; 1,944}
Roadnet-M 10,502 2,929 {1,145; 11,456; 114,561}
Roadnet-L 93,564 26,479 {24,649; 246,495; 2,464,950}
The statistics of the datasets used for the simulator performance comparison experiments are shown in Table A2. For hardware settings, all simulations are conducted in the same hardware environment with an Intel(R) Xeon(R) Platinum 8462Y CPU (64 threads) and an NVIDIA GeForce RTX 4090 GPU.
Appendix D Visualization of Simulator Realism Comparison
Figure A3: Visualization and comparison of road traffic status from (a) the Shenzhen dataset, (b) our proposed simulator, and (c) CityFlow. (best view in color)
Appendix E Datasets for Transportation System Optimization Benchmarking